### v2

@@ 4.1.2 Stratego
Unlike most board games, which are perfect information games, Stratego, a two-player imperfect information board game, distinguishes itself by incorporating elements of memory, deduction, and bluffing. This complexity is further amplified by the game’s long episode length and many potential game states, estimated $10^{535}$ [104]. The game is divided into two phases: the deployment phase, where players secretly arrange their units, and the game-play phase, where the objective is to deduce the opponent’s setup and capture their flag. The depth and computational complexity of the game remained a challenge until breakthroughs such as DeepNash [104] showed promising advances in AI’s ability to tackle it. DeepNash scales up evolution-theory-based self-play method R-NaD [105] (discussed in Sec. 3.3.5) to neural R-NaD. It employs a neural network with four heads: one for value prediction, one for the deployment phase, one for piece selection, and one for piece displacement. Neural Replicator Dynamics [52] is utilized to obtain the approximate fixed-point policy. DeepNash holds the third-place ranking among all professional Gravon Stratego players and wins nearly every match against existing Stratego bots.
@@ 4.2.1 Texas Hold’em
Heads-up limit Texas Hold’em (HULHE), the simplest form with approximately $3.16\times 10^{17}$ game states, was not solved until the introduction of Cepheus [10]. It utilizes fixed-point arithmetic with compression and regret-minimization-based self-play method CFR+ [122] to address the issues of storage and computation, respectively, resulting in superhuman performance. Deep CFR [14] combines regret-minimization-based self-play method CFR with deep neural networks. Furthermore, some studies do not adopt the regret-minimization-based series of self-play; instead, they use the traditional self-play method. NSFP [51] introduces a self-play method combined with end-to-end RL training, also achieving competitive performance in HULHE. Poker-CNN [140] utilizes self-play to learn convolutional networks to solve a video version of HULHE.
After solving HULHE, research focus shifts to heads-up no-limit Texas Hold’em (HUNL) with significantly larger game states, approximately $10^{164}$ [63]. Thus, traversing the game tree as Cepheus does in HULHE is impossible. DeepStack [95] is a regret-minimization-based self-play method combined with continual re-solving. This method focuses on a subtree of limited depth and breadth and estimates the outcomes of the furthest reaches of this subtree. Libratus [18] develops a blueprint strategy, leveraging an enhanced version of regret-minimization-based self-play method MCCFR [72]. Recursive Belief-based Learning (ReBeL) [12] introduces the public belief state (PBS), transforming the imperfect-information game into a perfect-information game with continuous state space. ReBeL utilizes the regret-minimization-based self-play method CFR-D, combined with RL and search, to train both the value and policy networks. Unlike the algorithms mentioned above, AlphaHoldem [149] does not use regret-minimization-based self-play. Instead, it proposes a $K$-best self-play method that always selects the top $K$ agents, based on their ELO ratings [36], to generate samples.
Pluribus [20], based on Libratus, addresses six-player no-limit Texas Hold’em. Similar to Libratus, Pluribus utilizes a regret-minimization-based self-play method MCCFR [72] to develop its blueprint strategy. It conducts a depth-limited search before executing actions. Different from Libratus, Pluribus maintains a streamlined policy pool of only four strategies, assuming that its opponents might adjust their strategies during gameplay among these four strategies, which allows Pluribus to manage complexity more efficiently.
@@ 4.2.2 DouDiZhu
DouDizhu (a.k.a. Fight the Landlord) is a three-player strategic card game. In this game, one player takes on the role of the landlord and competes against the other two players, the peasants. The game is played in two main stages: the bidding stage and the card play stage. During the bidding stage, players vie to become the landlord. During the card play stage, players take turns playing cards in various combinations, intending to be the first to empty their hands. DouDiZhu is characterized by imperfect information, as players can only see their own cards and the cards played. The essence of the game lies in cooperation among peasants and competition between the two factions. It has an estimated $10^{76}\sim 10^{106}$ possible game states [147] and an action space comprising $27472$ possible moves [141].
DeltaDou [61] is the first algorithm to achieve expert-level performance in DouDizhu. Similar to AlphaZero [116], DeltaDou utilizes vanilla self-play to generate samples to train policy and value networks. DouZero [148] also uses vanilla self-play to generate data. It reduces training costs by opting for a sampling method over the tree search approach. Based on DouZero, DouZero+ [150] incorporates opponent modeling, assisting the agent in making more informed decisions. PerfectDou [141] utilizes vanilla self-play to train a Perfect-Training-Imperfect-Execution framework.
@@ 4.2.3 Mahjong
Mahjong has evolved into various global variants, including the famous Japanese version known as Riichi Mahjong. This game is typically played by four players who must navigate both the visible aspects of the game, such as discarded tiles, and the hidden elements, like their own hand and the unseen public tiles. The strategic depth and complexity of Mahjong pose significant challenges. Despite ongoing research, these findings have yet to reach expert human levels [71, 40, 93]. Difficulties are navigating incomplete information, dynamically adapting strategies to multiple opponents, and contending with complex winning rules and an enormous number of possible game states, estimated at $10^{169}$ [147]. Suphx [77] is recognized as one of the first algorithms to master Mahjong, achieving a performance level comparable to expert human players, precisely 10 dan on Tenhou [127], the most popular online Mahjong platform. Initially, Suphx employs supervised learning, utilizing expert data to train its model. It then advances its capabilities through vanilla self-play combined with RL. Similarly, NAGA [128], developed by Dwango Media Village, and LuckyJ [125], designed by Tencent, have also achieved the rank of 10 dan on Tenhou. Furthermore, LuckyJ has even defeated human professional players.
@@ 4.3 Video Games
In contrast to traditional board games and card games, video games often feature real-time actions, long trajectories, and increased complexity stemming from a broader range of actions and observations. Here, we illustrate some representative video games to demonstrate how self-play has advanced AI within these games.
@@ 4.3.1 StarCraft II
StarCraft is a real-time strategy (RTS) game. It has three distinct species: the Terrans, Zerg, and Protoss, each with unique units and strategic options that enhance the gameplay complexity. Renowned for its balanced gameplay, strategic depth, and intense competitiveness, the game challenges players to gather resources, construct bases, and build armies. Victory requires meticulous planning and tactical execution, with defeat occurring when a player loses all their buildings. AlphaStar [129] dominates the 1v1 mode competitions in StarCraft II and has defeated professional players. Its framework is similar to AlphaGo [115], initially utilizing supervised learning to train the policy with expert data. Subsequently, it uses a hierarchical self-play method combined with end-to-end RL to train the networks. More specifically, The proposed self-play method divides all the agents into three types: main agents, league exploiters, and main exploiters. It maintains a policy pool of past players that records all these types of agents. Main agents engage in traditional self-play algorithms like FSP and PFSP, competing against main agents themselves and other agents in the policy pool. They are periodically added to the pool and never reset. League exploiters also use a traditional self-play method PFSP to play against all policy pool agents, added to the pool if they show a high win rate and potentially reset to expose global blind spots. Main exploiters only compete with main agents to improve their robustness, are added to the pool after achieving a high win rate or specific training steps, and are reset upon each addition. Among those three agent types, the main agent is the core agent and embodies the final AlphaStar strategy. However, the introduction of three agent types significantly increases the training computation. Further studies [48, 131, 57] have enhanced the league self-play training procedure.
@@ 4.3.2 MOBA Games
Multiplayer Online Battle Arena (MOBA) games are a popular video game genre that blends RTS with role-playing elements. In typical MOBA games, two teams of players control their unique characters, known as heroes, and compete to destroy the opposing team’s main structure, often referred to as the base. Each hero has distinct abilities and plays a specific role within the team, such as Warrior, Tank, or Support. Managing multiple lanes and battling under the fog of war, which obscures parts of the map, are critical aspects of the gameplay. Popular examples of MOBA games include League of Legends, Dota 2, and Honor of Kings.
OpenAI Five [7] defeated the world champion team in a simplified version of Dota 2 that featured a limited pool of heroes and certain banned items. The training process introduces a self-play method that combines two traditional self-play algorithms: with an 80% probability of engaging in vanilla self-play and a 20% probability of employing a technique similar to the PFSP used in AlphaStar [129]. This technique selects each policy from the policy pool based on its quality score, which is continuously updated from competition results throughout training. Higher quality scores increase the likelihood of a policy being selected. OpenAI Five also requires extensive training resources.
Another notable MOBA game, especially viral in China, is Honor of Kings. The 1v1 mode is conquered by [144], which boasts a significant winning rate against top professional players. It utilizes a traditional self-play method $\delta$-uniform self-play to train the RL networks. The 5v5 mode was later mastered by [143] as well. Unlike OpenAI Five, this work expands the hero pool to 40 heroes, exponentially increasing the possible combinations of lineups. It proposes a new self-play method referred to as curriculum self-play learning (CSPL). Specifically, the training process is divided into three stages. The first stage involves training fixed lineups through self-play and utilizing human data to balance the two teams to aid policy improvements. The second stage employs multi-teacher policy distillation to produce a distilled model. The final stage uses this distilled model as the initial model for another round of self-play with randomly picked lineups. This approach defeats professional player teams. The self-play generated data is also used to learn compelling lineup drafting by utilizing MCTS and neural networks [25].
@@ 4.3.3 Google Research Football
Google Research Football (GRF) [70] is an open-source football simulator emphasizing high-level actions. It initially offers two scenarios: the football benchmark and the football academy with 11 specific tasks. Here, we focus exclusively on the football benchmark because it presents a more complex scenario that better demonstrates the effects of self-play. GRF is particularly challenging due to the need for cooperation among teammates and competition against opposing teams. It features long trajectories with 3000 steps per round, stochastic transitions, and sparse rewards.
WeKick [153] claimed victory in the GRF competition on Kaggle [45], which simplifies the game dynamics by allowing competitors to control only one player, either the ball carrier on offense or the nearest defender on defense. WeKick employed self-play strategies similar to those used in league training [129]. It initializes its opponent policy pool using strategies developed through RL and Generative Adversarial Imitation Learning (GAIL) [55] to facilitate the training process.
Further research delves into the full football game rather than the simplified version. Team-PSRO [91] extends PSRO series of self-play algorithms to team games, outperforming baselines in the 4v4 version of the full GRF. In the context of the 11v11 version, where the goalkeeper is rule-based controlled, TiKick [58] utilizes vanilla self-play to collect samples and then employs imitation learning. Fictitious Cross-Play (FXP) [137] proposes a new self-play method similar to AlphaStar [129]. It introduces two populations: the main population and the counter population. Policies in the counter population improve solely by cross-playing with policies in the main population as opponents, while policies in the main population engage in playing with policies from both populations. FXP achieves a win rate of over 94% against TiKick. TiZero [78], a follow-up to TiKick, combines curriculum learning with FSP and PFSP [129] to avoid expert data reliance, achieving a higher TrueSkill rating [53] than TiKick.
@@
Board Game Chess 2 $\checkmark$ $10^{45}$ AlphaZero [116] Vanilla SP $\checkmark$ $\times$
MuZero [112] Vanilla SP $\checkmark$ $\times$
Go 2 $\checkmark$ $10^{360}$ AlphaGo [115] FSP $\checkmark$ $\checkmark$
AlphaGo Zero [117] Vanilla SP $\checkmark$ $\times$
AlphaZero [116] Vanilla SP $\checkmark$ $\times$
MuZero [112] Vanilla SP $\checkmark$ $\times$
Stratego 2 $\times$ $10^{535}$ DeepNash [104] R-NaD $\times$ $\times$
Card Game HULHE 2 $\times$ $10^{17}$ Cepheus [10] CFR+ $\checkmark$ $\times$
NFSP [51] NFSP $\times$ $\times$
HUNL 2 $\times$ $10^{164}$ DeepStack [95] CFR+ $\checkmark$ $\times$
Libratus [18] MCCFR $\checkmark$ $\times$
ReBel [12] CFR-D / FSP $\checkmark$ $\times$
Six-player No-limit Texas Hold’em 6 $\times$ $>10^{164}$ Pluribus [20] MCCFR $\checkmark$ $\times$
DouDiZhu 3 $\times$ $10^{76}\sim 10^{106}$ DeltaDou [61] Vanilla SP $\checkmark$ $\times$
DouZero [148] Vanilla SP $\times$ $\checkmark$
PerfectDou [141] Vanilla SP $\times$ $\times$
Mahjong 4 $\times$ $10^{169}$ Suphx [77] Vanilla SP $\checkmark$ $\checkmark$
Video Game StarCraft II 2 $\times$ / AlphaStar [129] FSP & PFSP $\times$ $\checkmark$
Dota 2 10 $\times$ OpenAI Five [7] Vanilla SP & PFSP $\times$ $\times$
Honor of Kings 10 $\times$ MOBA AI [143] $\delta$-uniform SP $\checkmark$ $\checkmark$
Google Research Football 22 $\checkmark$ TiZero [78] FSP & PFSP $\times$ $\times$

### v3

@@ 4.1.2 Stratego
Unlike most board games, which are perfect information games, Stratego, a two-player imperfect information board game, distinguishes itself by incorporating elements of memory, deduction, and bluffing. This complexity is further amplified by the game’s long episode length and many potential game states, estimated $10^{535}$ (Perolat et al., 2022). The game is divided into two phases: the deployment phase, where players secretly arrange their units, and the game-play phase, where the objective is to deduce the opponent’s setup and capture their flag. The depth and computational complexity of the game remained a challenge until breakthroughs such as DeepNash (Perolat et al., 2022) showed promising advances in AI’s ability to tackle it. DeepNash scales up evolution-theory-based self-play method R-NaD (Perolat et al., 2021) (discussed in Sec. 3.3.5) to neural R-NaD. It employs a neural network with four heads: one for value prediction, one for the deployment phase, one for piece selection, and one for piece displacement. Neural Replicator Dynamics (Hennes et al., 2020) is utilized to obtain the approximate fixed-point policy. DeepNash holds the third-place ranking among all professional Gravon Stratego players and wins nearly every match against existing Stratego bots.
@@ 4.2.1 Texas Hold’em
Heads-up limit Texas Hold’em (HULHE), the simplest form with approximately $3.16\times 10^{17}$ game states, was not solved until the introduction of Cepheus (Bowling et al., 2015). It utilizes fixed-point arithmetic with compression and regret-minimization-based self-play method CFR+ (Tammelin, 2014) to address the issues of storage and computation, respectively, resulting in superhuman performance. Deep CFR (Brown et al., 2019) combines regret-minimization-based self-play method CFR with deep neural networks. Furthermore, some studies do not adopt the regret-minimization-based series of self-play; instead, they use the traditional self-play method. NSFP (Heinrich & Silver, 2016) introduces a self-play method combined with end-to-end RL training, also achieving competitive performance in HULHE. Poker-CNN (Yakovenko et al., 2016) utilizes self-play to learn convolutional networks to solve a video version of HULHE.
After solving HULHE, research focus shifts to heads-up no-limit Texas Hold’em (HUNL) with significantly larger game states, approximately $10^{164}$ (Johanson, 2013). Thus, traversing the game tree as Cepheus does in HULHE is impossible. DeepStack (Moravčík et al., 2017) is a regret-minimization-based self-play method combined with continual re-solving. This method focuses on a subtree of limited depth and breadth and estimates the outcomes of the furthest reaches of this subtree. Libratus (Brown & Sandholm, 2018) develops a blueprint strategy, leveraging an enhanced version of regret-minimization-based self-play method MCCFR (Lanctot et al., 2009). Recursive Belief-based Learning (ReBeL) (Brown et al., 2020) introduces the public belief state (PBS), transforming the imperfect-information game into a perfect-information game with continuous state space. ReBeL utilizes the regret-minimization-based self-play method CFR-D, combined with RL and search, to train both the value and policy networks. Unlike the algorithms mentioned above, AlphaHoldem (Zhao et al., 2022a) does not use regret-minimization-based self-play. Instead, it proposes a $K$-best self-play method that always selects the top $K$ agents, based on their ELO ratings (Elo & Sloan, 1978), to generate samples.
Pluribus (Brown & Sandholm, 2019b), based on Libratus, addresses six-player no-limit Texas Hold’em. Similar to Libratus, Pluribus utilizes a regret-minimization-based self-play method MCCFR (Lanctot et al., 2009) to develop its blueprint strategy. It conducts a depth-limited search before executing actions. Different from Libratus, Pluribus maintains a streamlined policy pool of only four strategies, assuming that its opponents might adjust their strategies during gameplay among these four strategies, which allows Pluribus to manage complexity more efficiently.
@@ 4.2.2 DouDiZhu
DouDizhu (a.k.a. Fight the Landlord) is a three-player strategic card game. In this game, one player takes on the role of the landlord and competes against the other two players, the peasants. The game is played in two main stages: the bidding stage and the card play stage. During the bidding stage, players vie to become the landlord. During the card play stage, players take turns playing cards in various combinations, intending to be the first to empty their hands. DouDiZhu is characterized by imperfect information, as players can only see their own cards and the cards played. The essence of the game lies in cooperation among peasants and competition between the two factions. It has an estimated $10^{76}\sim 10^{106}$ possible game states (Zha et al., 2019) and an action space comprising $27472$ possible moves (Yang et al., 2022).
DeltaDou (Jiang et al., 2019) is the first algorithm to achieve expert-level performance in DouDizhu. Similar to AlphaZero (Silver et al., 2018), DeltaDou utilizes vanilla self-play to generate samples to train policy and value networks. DouZero (Zha et al., 2021) also uses vanilla self-play to generate data. It reduces training costs by opting for a sampling method over the tree search approach. Based on DouZero, DouZero+ (Zhao et al., 2022b) incorporates opponent modeling, assisting the agent in making more informed decisions. PerfectDou (Yang et al., 2022) utilizes vanilla self-play under a Perfect-Training-Imperfect-Execution framework.
@@ 4.2.3 Mahjong
Mahjong has evolved into various global variants, including the famous Japanese version known as Riichi Mahjong. This game is typically played by four players who must navigate both the visible aspects of the game, such as discarded tiles, and the hidden elements, like their own hand and the unseen public tiles. The strategic depth and complexity of Mahjong pose significant challenges. Despite ongoing research, these findings have yet to reach expert human levels (Kurita & Hoki, 2020; Gao et al., 2019; Mizukami & Tsuruoka, 2015). Difficulties are navigating incomplete information, dynamically adapting strategies to multiple opponents, and contending with complex winning rules and an enormous number of possible game states, estimated at $10^{169}$ (Zha et al., 2019). Suphx (Li et al., 2020) is recognized as one of the first algorithms to master Mahjong, achieving a performance level comparable to expert human players, precisely 10 dan on Tenhou (Tsunoda, ), the most popular online Mahjong platform. Initially, Suphx employs supervised learning, utilizing expert data to train its model. It then advances its capabilities through vanilla self-play combined with RL. Similarly, NAGA (Village, ), developed by Dwango Media Village, and LuckyJ (Tencent, ), designed by Tencent, have also achieved the rank of 10 dan on Tenhou. Furthermore, LuckyJ has even defeated human professional players.
@@ 4.3 Video Games
In contrast to traditional board games and card games, video games often feature real-time actions, long trajectories, and increased complexity stemming from a broader range of actions and observations. We illustrate representative video games showcasing self-play’s impact.
@@ 4.3.1 StarCraft II
StarCraft is a real-time strategy (RTS) game. It has three distinct species: the Terrans, Zerg, and Protoss, each with unique units and strategic options that enhance the gameplay complexity. Renowned for its balanced gameplay, strategic depth, and intense competitiveness, the game challenges players to gather resources, construct bases, and build armies. Victory requires meticulous planning and tactical execution, with defeat occurring when a player loses all their buildings. AlphaStar (Vinyals et al., 2019) dominates the 1v1 mode competitions in StarCraft II and has defeated professional players. Its framework is similar to AlphaGo (Silver et al., 2016), initially utilizing supervised learning to train the policy with expert data. Subsequently, it uses a hierarchical self-play method combined with end-to-end RL to train the networks. More specifically, The proposed self-play method divides all the agents into three types: main agents, league exploiters, and main exploiters. It maintains a policy pool of past players that records all these types of agents. Main agents engage in traditional self-play algorithms like FSP and PFSP, competing against main agents themselves and other agents in the policy pool. They are periodically added to the pool and never reset. League exploiters also use a traditional self-play method PFSP to play against all policy pool agents, added to the pool if they show a high win rate and potentially reset to expose global blind spots. Main exploiters only compete with main agents to improve their robustness, are added to the pool after achieving a high win rate or specific training steps, and are reset upon each addition. Among those three agent types, the main agent is the core agent and embodies the final AlphaStar strategy. However, the introduction of three agent types significantly increases the training computation. Further studies (Han et al., 2020; Wang et al., 2021; Huang et al., 2024) have enhanced the league self-play training procedure.
@@ 4.3.2 MOBA Games
Multiplayer Online Battle Arena (MOBA) games are a popular video game genre that blends RTS with role-playing elements. In typical MOBA games, two teams of players control their unique characters, known as heroes, and compete to destroy the opposing team’s main structure, often referred to as the base. Each hero has distinct abilities and plays a specific role within the team, such as Warrior, Tank, or Support. Managing multiple lanes and battling under the fog of war, which obscures parts of the map, are critical aspects of the gameplay.
OpenAI Five (Berner et al., 2019) defeated the world champion team in a simplified version of Dota 2 that featured a limited pool of heroes and certain banned items. The training process introduces a self-play method that combines two traditional self-play algorithms: with an 80% probability of engaging in vanilla self-play and a 20% probability of employing a technique similar to the PFSP used in AlphaStar (Vinyals et al., 2019). This technique selects each policy from the policy pool based on its quality score, which is continuously updated from competition results throughout training. Higher quality scores increase the likelihood of a policy being selected. OpenAI Five also requires extensive training resources.
Another notable MOBA game, especially viral in China, is Honor of Kings. The 1v1 mode is conquered by Ye et al. (2020b), which boasts a significant winning rate against top professional players. It utilizes a traditional self-play method $\delta$-uniform self-play to train the RL networks. The 5v5 mode was later mastered by Ye et al. (2020a) as well. Unlike OpenAI Five, this work expands the hero pool to 40 heroes, exponentially increasing the possible combinations of lineups. It proposes a new self-play method referred to as curriculum self-play learning (CSPL). Specifically, the training process is divided into three stages. The first stage involves training fixed lineups through self-play and utilizing human data to balance the two teams to aid policy improvements. The second stage employs multi-teacher policy distillation to produce a distilled model. The final stage uses this distilled model as the initial model for another round of self-play with randomly picked lineups. This approach defeats professional player teams. The self-play generated data is also used to learn compelling lineup drafting by utilizing MCTS and neural networks (Chen et al., 2021).
@@ 4.3.3 Google Research Football
Google Research Football (GRF) (Kurach et al., 2020) is an open-source football simulator emphasizing high-level actions. It initially offers two scenarios: the football benchmark and the football academy with 11 specific tasks. Here, we focus exclusively on the football benchmark because it presents a more complex scenario that better demonstrates the effects of self-play. GRF is particularly challenging due to the need for cooperation among teammates and competition against opposing teams. It features long trajectories with 3000 steps per round, stochastic transitions, and sparse rewards.
WeKick (Ziyang Li, 2020) claimed victory in the GRF competition on Kaggle (Google, 2020), which simplifies the game dynamics by allowing competitors to control only one player, either the ball carrier on offense or the nearest defender on defense. It employed self-play strategies similar to those used in league training (Vinyals et al., 2019). It initializes its opponent policy pool using strategies developed through RL and Generative Adversarial Imitation Learning (GAIL) (Ho & Ermon, 2016) to facilitate the training process.
Further research delves into the full football game rather than the simplified version. Team-PSRO (McAleer et al., 2023) extends PSRO series of self-play algorithms to team games, outperforming baselines in the 4v4 version of the full GRF. In the context of the 11v11 version, where the goalkeeper is rule-based controlled, TiKick (Huang et al., 2021a) utilizes vanilla self-play to collect samples and then employs imitation learning. Fictitious Cross-Play (FXP) (Xu et al., 2023) proposes a new self-play method similar to AlphaStar (Vinyals et al., 2019). It introduces two populations: the main population and the counter population. Policies in the counter population improve solely by cross-playing with policies in the main population as opponents, while policies in the main population engage in playing with policies from both populations. FXP achieves a win rate of over 94% against TiKick. TiZero (Lin et al., 2023), a follow-up to TiKick, combines curriculum learning with FSP and PFSP (Vinyals et al., 2019) to avoid expert data reliance, achieving a higher TrueSkill rating (Herbrich et al., 2006) than TiKick.
@@
Board Game Chess 2 $\checkmark$ $10^{45}$ AlphaZero Vanilla SP $\checkmark$ $\times$
MuZero Vanilla SP $\checkmark$ $\times$
Go 2 $\checkmark$ $10^{360}$ AlphaGo FSP $\checkmark$ $\checkmark$
AlphaGo Zero Vanilla SP $\checkmark$ $\times$
AlphaZero Vanilla SP $\checkmark$ $\times$
MuZero Vanilla SP $\checkmark$ $\times$
Stratego 2 $\times$ $10^{535}$ DeepNash R-NaD $\times$ $\times$
Card Game HULHE 2 $\times$ $10^{17}$ Cepheus CFR+ $\checkmark$ $\times$
NFSP NFSP $\times$ $\times$
HUNL 2 $\times$ $10^{164}$ DeepStack CFR+ $\checkmark$ $\times$
Libratus MCCFR $\checkmark$ $\times$
ReBel CFR-D / FSP $\checkmark$ $\times$
Six-player No-limit Texas Hold’em 6 $\times$ $>10^{164}$ Pluribus MCCFR $\checkmark$ $\times$
DouDiZhu 3 $\times$ $10^{76}\sim 10^{106}$ DeltaDou Vanilla SP $\checkmark$ $\times$
DouZero Vanilla SP $\times$ $\checkmark$
PerfectDou Vanilla SP $\times$ $\times$
Mahjong 4 $\times$ $10^{169}$ Suphx Vanilla SP $\checkmark$ $\checkmark$
Video Game StarCraft II 2 $\times$ / AlphaStar FSP & PFSP $\times$ $\checkmark$
Dota 2 10 $\times$ OpenAI Five Vanilla SP & PFSP $\times$ $\times$
Honor of Kings 10 $\times$ MOBA AI $\delta$-uniform SP $\checkmark$ $\checkmark$
Google Research Football 22 $\checkmark$ TiZero FSP & PFSP $\times$ $\times$